The base of an isosceles triangle is two-fifths the length of each of the congruent sides. The perimeter of the triangle is 60 cm. How long is the base
Answer by jean14(9)

perimeter is simply the sum of all sides.
Congruent sides have same length.
solve;
Let x be the length of the 2 congruent sides
2/5 x be the base length

Make equation when perimeter is 60, solve for x.
60 = x + x + 2x/5
60 = 2x + 2x/5
60 = 2 2x/5
60=12x/5 multiply both sides by 5
5(60) = (12x/5) 5
300 =12x divide both sides by 12
300/12 = 12x/12
25 = x
The value of each congruent sides is 25cm. And the base is 2/5(25)=base
Then,, 50/5=10cm is the base length...

Check:
60cm = 25cm +25cm + 2/5(25cm)
60cm = 50cm + 10cm
60cm =60cm True